# WaveC2R: Wavelet-Driven Coarse-to-Refined Hierarchical Learning for Radar Retrieval

### v1

@@ Abstract
Satellite-based radar retrieval methods are widely employed to fill coverage gaps in ground-based radar systems, especially in remote areas affected by terrain blockage and limited detection range. Existing methods predominantly rely on overly simplistic spatial-domain architectures constructed from a single data source, limiting their ability to accurately capture complex precipitation patterns and sharply defined meteorological boundaries. To address these limitations, we propose WaveC2R, a novel wavelet-driven coarse-to-refined framework for radar retrieval. WaveC2R integrates complementary multi-source data and leverages frequency-domain decomposition to separately model low-frequency components for capturing precipitation patterns and high-frequency components for delineating sharply defined meteorological boundaries. Specifically, WaveC2R consists of two stages (i) Intensity-Boundary Decoupled Learning, which leverages wavelet decomposition and frequency-specific loss functions to separately optimize low-frequency intensity and high-frequency boundaries; and (ii) Detail-Enhanced Diffusion Refinement, which employs frequency-aware conditional priors and multi-source data to progressively enhance fine-scale precipitation structures while preserving coarse-scale meteorological consistency. Experimental results on the publicly available SEVIR dataset demonstrate that WaveC2R achieves state-of-the-art performance in satellite-based radar retrieval, particularly excelling at preserving high-intensity precipitation features and sharply defined meteorological boundaries. Our project is available at https://spring-lovely.github.io/WaveC2R/.
@@ 1 Introduction
Satellite-based radar retrieval plays a critical role in filling coverage gaps of ground-based radar systems, providing essential precipitation monitoring in data-sparse regions (Park et al. 2025; Wen et al. 2024). Geostationary satellites such as Himawari-8, GOES-R, and Fengyun-4A provide high-resolution meteorological observations (Liu et al. 2020; Jing et al. 2022) for radar composite reflectivity reconstruction. These satellites capture visible, infrared, and lightning data, which collectively enhance characterization of complex precipitation patterns in regions lacking ground-based radar coverage (Li et al. 2023; Yang et al. 2023a).
@@
Traditional retrieval methods, such as radiative transfer models (RTMs) (Di et al. 2018) and variational methods (Yin et al. 2021), rely on physical models and statistical techniques but are computationally intensive with limited accuracy. Deep learning methods have emerged as promising alternatives due to their superior data processing capabilities (Jiang et al. 2024; Park et al. 2025). However, current deep learning-based approaches still face several critical limitations (Fig.1).
@@
Single-source data limitations. Most current approaches rely solely on single-source satellite data, typically infrared channels. This neglects complementary information from alternative data sources. Visible observations provide crucial cloud-top morphological details during daylight hours that enhance structural identification. Similarly, lightning data serves as a direct indicator of convective intensity and enables precise storm core localization (Yu et al. 2025; Zheng et al. 2024; Zhao et al. 2024b; Chen et al. 2024). Without these multi-source inputs, current methods struggle with incomplete scene representation, limiting their ability to accurately characterize severe convective cells (Jin, Lin, and Xu 2023; Niu et al. 2024).
@@
Frequency domain optimization challenges. Accurate characterization of convective-scale structures requires effective modeling of meteorological systems’ multi-frequency characteristics (Jin, Lin, and Xu 2023). However, existing satellite-based radar retrieval methods predominantly employ spatial domain optimization with simplistic architectures and L2 loss functions (Zhao et al. 2024a). This approach fails to capture the inherent frequency characteristics of weather phenomena, resulting in oversmoothed gradients, blurred precipitation boundaries, and underestimated high-intensity convective cores (Yang et al. 2023b; Hilburn, Ebert-Uphoff, and Miller 2020).
@@
Limited diffusion model exploration in radar retrival. Diffusion models offer significant potential for satellite-based radar retrieval through their ability to model complex probability distributions and progressively refine meteorological structures. Despite proven success in various image generation tasks, diffusion models remain unexplored in satellite-based radar retrieval. Traditional approaches struggle with recovering high-frequency meteorological details, particularly in convective regions with steep gradients (Gong et al. 2024; Wen et al. 2024; Li et al. 2024).
@@ 2.1 Radar Retrieval Using Satellite Observations
Deep learning has emerged as a transformative approach for radar composite reflectivity (CREF) reconstruction by leveraging satellite observations. UNet-based models have become predominant due to their encoder-decoder structure and multi-scale feature extraction capabilities, with substantial improvements demonstrated through multi-band satellite integration (Jin, Lin, and Xu 2023; Hilburn, Ebert-Uphoff, and Miller 2020; Zhao et al. 2024a). Alternative architectural innovations have emerged to address UNet limitations, including residual deep forest models (Bao et al. 2024) and integrated frameworks combining multi-band infrared data (Wang et al. 2020). Recent investigations have highlighted the importance of integrating diverse satellite observations encompassing visible, infrared, and lightning data to capture spatiotemporal information (Si et al. 2023; Yu et al. 2025; Zheng et al. 2024; Zhao et al. 2024b; Chen et al. 2024). However, existing methodologies continue to face challenges in effectively integrating multi-channel data, often relying on simple concatenation strategies that fail to exploit the complementary strengths of different satellite observations.
@@ 2.2 Frequency-Domain and Wavelet Transform in Meteorological Applications
Wavelet transform has established itself as a powerful technique for multi-scale meteorological analysis, offering distinctive advantages over traditional Fourier methods when processing non-stationary weather signals (Saoud, Al-Marzouqi, and Deriche 2021). The integration of frequency-domain techniques with deep learning has yielded promising results, with improvements achieved through Fast Fourier Transform integration in weather forecasting frameworks (Li et al. 2025) and specialized frequency-domain learning methods (Wang et al. 2024). The multi-resolution capabilities of wavelet transforms have proven particularly valuable for precipitation systems, with enhanced predictions resulting from combining wavelet transforms with time series models (Ahn and Hur 2023; Mammedov, Olugu, and Farah 2022) and frequency domain feature extraction fusion mechanisms (Liao, Gao, and Li 2024). However, existing wavelet-based approaches in meteorology primarily focus on mathematical decomposition without exploring the distinct physical roles of different frequency components in precipitation systems.
@@ 3.4 Stage I: Intensity-Boundary Decoupled Learning
Traditional spatial-domain losses treat all frequency components equally, leading to energy imbalance where low-frequency intensity patterns dominate optimization while high-frequency boundary details are under-represented. Although Fourier Global Loss (FGL)(Yan et al. 2024) effectively captures global motion trends, it struggles with local positioning accuracy due to its global frequency spectrum operation, making it insufficient for capturing localized high-frequency structures that define precipitation boundaries.
@@ 3.5 Stage II: Detail-Enhanced Diffusion Refinement
Pred. AA-TransUnet (Yang 2022) MSE 0.302 0.437 0.328 0.407 0.318
@@
Smaat-Unet (Kevin 2021) MSE 0.293 0.421 0.309 0.364 0.362
@@
Gen. Diffcast (Yu et al. 2024) 0.310 0.448 0.391 0.579 0.236
UVCGAN (Torbunov et al. 2023) 0.089 0.131 0.125 0.254 0.362
Pix2pix (Akter 2024) 0.240 0.355 0.266 0.337 0.349
MeanFlow (Geng et al. 2025) 0.200 0.300 0.386 0.583 0.300
@@
Pred. AA-TransUnet (Yang 2022) MSE 0.506 0.329 0.317 0.262 0.095 0.638 0.483 0.476 0.412 0.174
@@
Smaat-Unet (Kevin 2021) MSE 0.516 0.324 0.313 0.256 0.055 0.646 0.478 0.471 0.404 0.105
@@
Gen. Diffcast (Yu et al. 2024) 0.501 0.322 0.321 0.282 0.126 0.629 0.473 0.479 0.436 0.223
UVCGAN (Torbunov et al. 2023) 0.228 0.094 0.063 0.043 0.015 0.306 0.144 0.104 0.074 0.028
Pix2pix (Akter 2024) 0.418 0.259 0.269 0.241 0.014 0.547 0.397 0.417 0.384 0.028
MeanFlow (Geng et al. 2025) 0.425 0.215 0.177 0.141 0.044 0.550 0.333 0.290 0.242 0.084
@@ 4.1 Implementation Details
Dataset. We utilize the SEVIR dataset (Veillette, Samsi, and Mattioli 2020), which integrates multi-source observations including visible data, infrared bands (IR069, IR107), vertically integrated liquid (VIL), and lightning data. The dataset contains over 20,000 storm events from 2017-2019, each with 4-hour temporal coverage at 1 km spatial resolution. We evaluate performance at VIL thresholds of 74, 133, 160, 181, and 219 kg/m². Training details are provided in Appendix B.1.
Evaluation Metrics. To ensure fair comparison with existing methods, we adopt the same evaluation metrics as FACL (Yan et al. 2024). Meteorological performance is assessed using Critical Success Index (CSI) and Heidke Skill Score (HSS) at multiple reflectivity thresholds, which are standard metrics in weather radar evaluation (Feng et al. 2025). Image quality evaluation uses Structural Similarity Index (SSIM). For perceptual assessment, we employ Learned Perceptual Image Patch Similarity (LPIPS).
Baseline Comparisons. We compare WaveC2R against CNN-Transformer hybrid models (AA-TransUnet (Yang 2022)), meteorology-oriented architectures (Earthformer (Gao et al. 2022)), CNN-based networks (Smaat-Unet (Kevin 2021)), and generative models including diffusion-based (Diffcast (Yu et al. 2024)), GAN-based methods (UVCGAN (Torbunov et al. 2023), Pix2Pix (Akter 2024)) and Flow-based methods (MeanFlow (Geng et al. 2025)). MeanFlow details are provided in Appendix A.2.
@@ 4.3 Ablation Studies
FIBL Loss Analysis. Comparing FIBL against MSE and FACL losses on Earthformer shows FIBL achieves the highest extreme-threshold performance (Table 2): CSI ($\gamma\!\geq\!219$) improves from 0.069 (MSE) to 0.149 (+116%), while maintaining superior boundary sharpness without artifacts. Fig. 6 visualizes FIBL’s effectiveness in preserving convective cores across diverse cases.

### v2

@@ Abstract
Satellite-based radar retrieval methods are widely employed to fill coverage gaps in ground-based radar systems, especially in remote areas affected by terrain blockage and limited detection range. Existing methods predominantly rely on overly simplistic spatial-domain architectures constructed from a single data source, limiting their ability to accurately capture complex precipitation patterns and sharply defined meteorological boundaries. To address these limitations, we propose WaveC2R, a novel wavelet-driven coarse-to-refined framework for radar retrieval. WaveC2R integrates complementary multi-source data and leverages frequency-domain decomposition to separately model low-frequency components for capturing precipitation patterns and high-frequency components for delineating sharply defined meteorological boundaries. Specifically, WaveC2R consists of two stages (i) Intensity-Boundary Decoupled Learning, which leverages wavelet decomposition and frequency-specific loss functions to separately optimize low-frequency intensity and high-frequency boundaries; and (ii) Detail-Enhanced Diffusion Refinement, which employs frequency-aware conditional priors and multi-source data to progressively enhance fine-scale precipitation structures while preserving coarse-scale meteorological consistency. Experimental results on the publicly available SEVIR dataset demonstrate that WaveC2R achieves state-of-the-art performance in satellite-based radar retrieval, particularly excelling at preserving high-intensity precipitation features and sharply defined meteorological boundaries.
Code–https://spring-lovely.github.io/WaveC2R/˝
@@ 1 Introduction
Satellite-based radar retrieval plays a critical role in filling coverage gaps of ground-based radar systems, providing essential precipitation monitoring in data-sparse regions (park2025data; wen2024duocast). Geostationary satellites such as Himawari-8, GOES-R, and Fengyun-4A provide high-resolution meteorological observations (liu2020noaa; jing2022cascaded) for radar composite reflectivity reconstruction. These satellites capture visible, infrared, and lightning data, which collectively enhance characterization of complex precipitation patterns in regions lacking ground-based radar coverage (li2023toward; yang2023multitask).
@@
Traditional retrieval methods, such as radiative transfer models (RTMs) (di2018enhancing) and variational methods (yin2021impact), rely heavily on physical models and statistical techniques. These methods are computationally intensive and often constrained in accuracy due to various factors. In recent years, deep learning methods have become a hot topic in radar retrieval research due to their powerful data processing capabilities and efficient predictive performance (jiang2024retrieving; park2025data). However, despite the notable progress in deep learning-based satellite radar retrieval, current methods still suffer from several critical limitations that restrict their performance (Fig.1).
@@
Single-source data limitations. Most current approaches rely solely on single-source satellite data, typically infrared channels. This neglects complementary information from alternative data sources. Visible observations provide crucial cloud-top morphological details during daylight hours that enhance structural identification. Similarly, lightning data serves as a direct indicator of convective intensity and enables precise storm core localization (yu2025integrating; zheng2024cross; zhao2024weathergfm; chen2024terra). Without these multi-source inputs, current methods struggle with incomplete scene representation, limiting their ability to accurately characterize severe convective cells (jin2023rda; niu2024fsrgan).
@@
Frequency domain optimization challenges. Accurate characterization of convective-scale structures requires effective modeling of meteorological systems’ multi-frequency characteristics (jin2023rda). However, existing satellite-based radar retrieval methods predominantly employ spatial domain optimization with simplistic architectures and L2 loss functions (zhao2024intelligent). This approach fails to capture the inherent frequency characteristics of weather phenomena, resulting in oversmoothed gradients, blurred precipitation boundaries, and underestimated high-intensity convective cores (yang2023deep; hilburn2020development).
@@
Limited diffusion model exploration in radar retrival. Diffusion models offer significant potential for satellite-based radar retrieval through their ability to model complex probability distributions and progressively refine meteorological structures. Despite proven success in various image generation tasks, diffusion models remain unexplored in satellite-based radar retrieval. Traditional approaches struggle with recovering high-frequency meteorological details, particularly in convective regions with steep gradients (gong2024cascast; wen2024duocast; li2024scrd).
@@ 2.1 Radar Retrieval Using Satellite Observations
Deep learning has emerged as a transformative approach for radar composite reflectivity (CREF) reconstruction by leveraging satellite observations. UNet-based models have become predominant due to their encoder-decoder structure and multi-scale feature extraction capabilities, with substantial improvements demonstrated through multi-band satellite integration (jin2023rda; hilburn2020development; zhao2024intelligent). Alternative architectural innovations have emerged to address UNet limitations, including residual deep forest models (bao2024infrared) and integrated frameworks combining multi-band infrared data (wang2020infrared). Recent investigations have highlighted the importance of integrating diverse satellite observations encompassing visible, infrared, and lightning data to capture spatiotemporal information (si2023novel; yu2025integrating; zheng2024cross; zhao2024weathergfm; chen2024terra). However, existing methodologies continue to face challenges in effectively integrating multi-channel data, often relying on simple concatenation strategies that fail to exploit the complementary strengths of different satellite observations.
@@ 2.2 Frequency-Domain and Wavelet Transform in Meteorological Applications
Wavelet transform has established itself as a powerful technique for multi-scale meteorological analysis, offering distinctive advantages over traditional Fourier methods when processing non-stationary weather signals (saoud2021wind). The integration of frequency-domain techniques with deep learning has yielded promising results, with improvements achieved through Fast Fourier Transform integration in weather forecasting frameworks (li2025climatellm) and specialized frequency-domain learning methods (wang2024fredf). The multi-resolution capabilities of wavelet transforms have proven particularly valuable for precipitation systems, with enhanced predictions resulting from combining wavelet transforms with time series models (ahn2023short; mammedov2022weather) and frequency domain feature extraction fusion mechanisms (liao2024wind). However, existing wavelet-based approaches in meteorology primarily focus on mathematical decomposition without exploring the distinct physical roles of different frequency components in precipitation systems.
@@ 3.4 Stage I: Intensity-Boundary Decoupled Learning
Traditional spatial-domain losses treat all frequency components equally, leading to energy imbalance where low-frequency intensity patterns dominate optimization while high-frequency boundary details are under-represented. Although Fourier Global Loss (FGL)(yan2024fourier) effectively captures global motion trends, it struggles with local positioning accuracy due to its global frequency spectrum operation, making it insufficient for capturing localized high-frequency structures that define precipitation boundaries.
@@ 3.5 Stage II: Detail-Enhanced Diffusion Refinement
Pred. AA-TransUnet (yang2022aa) MSE 0.302 0.437 0.328 0.407 0.318
@@
Smaat-Unet (trebing2021smaat) MSE 0.293 0.421 0.309 0.364 0.362
@@
Gen. Diffcast (yu2024diffcast) 0.310 0.448 0.391 0.579 0.236
UVCGAN (torbunov2023uvcgan) 0.089 0.131 0.125 0.254 0.362
Pix2pix (akter2024generative) 0.240 0.355 0.266 0.337 0.349
MeanFlow (geng2025mean) 0.200 0.300 0.386 0.583 0.300
@@
Pred. AA-TransUnet (yang2022aa) MSE 0.506 0.329 0.317 0.262 0.095 0.638 0.483 0.476 0.412 0.174
@@
Smaat-Unet (trebing2021smaat) MSE 0.516 0.324 0.313 0.256 0.055 0.646 0.478 0.471 0.404 0.105
@@
Gen. Diffcast (yu2024diffcast) 0.501 0.322 0.321 0.282 0.126 0.629 0.473 0.479 0.436 0.223
UVCGAN (torbunov2023uvcgan) 0.228 0.094 0.063 0.043 0.015 0.306 0.144 0.104 0.074 0.028
Pix2pix (akter2024generative) 0.418 0.259 0.269 0.241 0.014 0.547 0.397 0.417 0.384 0.028
MeanFlow (geng2025mean) 0.425 0.215 0.177 0.141 0.044 0.550 0.333 0.290 0.242 0.084
@@ 4.1 Implementation Details
Dataset. We utilize the SEVIR dataset (veillette2020sevir), which integrates multi-source observations including visible data, infrared bands (IR069, IR107), vertically integrated liquid (VIL), and lightning data. The dataset contains over 20,000 storm events from 2017-2019, each with 4-hour temporal coverage at 1 km spatial resolution. We evaluate performance at VIL thresholds of 74, 133, 160, 181, and 219 kg/m². Training details are provided in Appendix B.1.
Evaluation Metrics. To ensure fair comparison with existing methods, we adopt the same evaluation metrics as FACL (yan2024fourier). Meteorological performance is assessed using Critical Success Index (CSI) and Heidke Skill Score (HSS) at multiple reflectivity thresholds, which are standard metrics in weather radar evaluation (fengperceptually). Image quality evaluation uses Structural Similarity Index (SSIM). For perceptual assessment, we employ Learned Perceptual Image Patch Similarity (LPIPS).
Baseline Comparisons. We compare WaveC2R against CNN-Transformer hybrid models (AA-TransUnet (yang2022aa)), meteorology-oriented architectures (Earthformer (gao2022earthformer)), CNN-based networks (Smaat-Unet (trebing2021smaat)), and generative models including diffusion-based (Diffcast (yu2024diffcast)), GAN-based methods (UVCGAN (torbunov2023uvcgan), Pix2Pix (akter2024generative)) and Flow-based methods (MeanFlow (geng2025mean)). MeanFlow details are provided in Appendix A.2.
@@ 4.3 Ablation Studies
FIBL Loss Analysis. Comparing FIBL against MSE and FACL losses on Earthformer shows FIBL achieves the highest extreme-threshold performance (see Table 2): CSI ($\gamma\!\geq\!219$) improves from 0.069 (MSE) to 0.149 (+116%), while maintaining superior boundary sharpness without artifacts. Fig. 6 visualizes FIBL’s effectiveness in preserving convective cores across diverse cases.